OpenStudy (anonymous):

a trapezoid has an area of 166.5in^2height 9, and one base 15in. what is the length of the other base?

OpenStudy (anonymous):

A=.5(b1+b2)*h correct

OpenStudy (anonymous):

and your givens are h=9 b1=15 A=166.5 in^2

OpenStudy (anonymous):

yea is their some kind of equation to find the other base?

OpenStudy (anonymous):

so if you just replace the formula with your given data you'll get 166.5=.5(15+b2)*9

OpenStudy (anonymous):

you can solve it with the regular equation since there is only one missing variable

OpenStudy (anonymous):

now snce everything on the right is multiplied by each other, you can get rid of the .5 by multiplying it with 9 to get 4.5

OpenStudy (anonymous):

so now you should have 166.5=4.5(15+b2)

OpenStudy (anonymous):

if you multiply this out you get\[166.5=67.5+4.5b_2\]

OpenStudy (anonymous):

you distribut it/

OpenStudy (anonymous):

yes distribute and then it becomes a simple problem to solve

OpenStudy (anonymous):

\[166.5-67.5=67.5-67.5+4.5b_2\]

OpenStudy (anonymous):

\[99=4.5b_2\]

OpenStudy (anonymous):

22?

OpenStudy (anonymous):

\[\frac{99}{4.5}=\frac{4.5b_2}{4.5}\]

OpenStudy (anonymous):

b2 = 22

OpenStudy (anonymous):

now to check use the original formula for A

OpenStudy (anonymous):

ok thanks ! but i still have a question for the other one when i got the answer that is not one of the answer choices?

OpenStudy (anonymous):

\[166.5=.5(15+22)*9=4.5(15+22)=4.5(37)=166.5\]

OpenStudy (anonymous):

what were the answer choices

OpenStudy (anonymous):

36squareroute3cm^2 44squareroute3cm^2 65cm^2 88cm^2

OpenStudy (anonymous):

65 is probably the answer since there i probably a few rounding errors let me just make sure

OpenStudy (anonymous):

ok that is what i thought too

OpenStudy (anonymous):

yeah if you do the triangle seperately and then the rectangle it comes out to be 64.495 cm^2

OpenStudy (anonymous):

if you use the trapezoid formula it comes out to 62.4